# To FP8 and Back Again: Quantifying the Effects of Reducing Precision on LLM Training Stability

### v1

@@ Abstract
The massive computational costs associated with large language model (LLM) pretraining have spurred great interest in reduced-precision floating-point representations to accelerate the process. As a result, the BrainFloat16 (BF16) precision has become the de facto standard for LLM training, with hardware support included in recent accelerators. This trend has gone even further in the latest processors, where FP8 has recently been introduced. However, prior experience with FP16, which was found to be less stable than BF16, raises concerns as to whether FP8, with even fewer bits than FP16, can be a cost-effective option for LLM training. We argue that reduced-precision training schemes must have similar training stability and hyperparameter sensitivities to their higher-precision counterparts in order to be cost-effective. However, we find that currently available methods for FP8 training are not robust enough to allow their use as economical replacements. This prompts us to investigate the stability of reduced-precision LLM training in terms of robustness across random seeds and learning rates. To this end, we propose new evaluation techniques and a new metric for quantifying loss landscape sharpness in autoregressive language models. By simulating incremental bit reductions in floating-point representations, we analyze the relationship between representational power and training stability with the intent of aiding future research into the field.
@@ 1 Introduction
Conversational large language models (LLMs), such as ChatGPT [1], Gemini [2, 3], Claude [4], Llama [5, 6], and HyperCLOVA [7], have captured the imagination of both academics and the public at large with their ability to communicate fluently with humans in natural language. However, these models require unprecedented amounts of computation to train, which has engendered interest in methods to improve training efficiency.
A popular method of improving computational performance is to reduce the bit count of the floating-point representations used for training [8, 9, 10]. This is because reading memory is the main bottleneck in modern processors, a problem known as the “memory wall” [11, 12]. Reducing the number of bits that each floating-point number uses can, therefore, accelerate the computation in proportion to the amount of memory reduced. For example, in processors that support it, computations in BrainFloat16 (BF16) [13] can have double the maximum throughput of single precision FP32. Furthermore, the FP32 data type, the highest precision data type used in deep learning, has only half the bits of FP64, the most widely used floating-point data type in scientific computing. The current best practice for LLM training is to use BF16 for most of the LLM training computation, with some sensitive portions, such as layer normalization [14], carried out in FP32.
As a natural extension of this development, 8-bit floating-point (FP8) [8, 15, 16, 10] and even 4-bit floating-point [9] data formats have been proposed to accelerate training even further. However, the naïve application of FP8 to LLM training is unstable and requires additional techniques to become viable. While several methods have been proposed to stabilize training LLMs with FP8, relatively little attention has been paid to quantifying the decrease in training stability compared to the standard mixed-precision BF16 training.
As the motivation behind the use of FP8 and other reduced-precision training schemes is cost reduction, our concern here is not whether LLM pretraining with FP8 is possible but whether it is profitable. For cost savings to be realized, the time per training step must be reduced while the number of training steps is kept to a similar number. Training stability is thus a crucial factor for cost-effective LLM training, considering that additional hyperparameter searches and restarts from training failures can outweigh any gains from the acceleration in raw computation. Therefore, for the newly proposed reduced-precision training schemes to be economical, the models trained on them must be similarly robust to hyperparameter choice and stochastic noise as models trained using higher precision.
@@
We motivate our line of inquiry with some surprising findings from experiments on the nanoGPT [17] codebase, an open-source implementation of GPT-2 pretraining, where we found that even the current best practice of mixed-precision BF16 can introduce training instabilities. When we compared BF16 and TensorFloat32 (TF32) runs, where we ran training for 5% of the dataset for each run, we found that the BF16 models diverged for 18 of 188 runs, or approximately 10% of all cases, despite using the same configurations as the default run. In contrast, no cases of loss divergence were found for the 70 TF32 models that were trained using different random seeds. We compare against the TF32 data type because NVIDIA GPUs do not offer tensor cores in FP32.
This is a surprising finding in light of the fact that most recent LLMs are trained with mixed precision BF16 without a comparison with training on TF32, which has three additional mantissa bits. However, a loss divergence rate of approximately 10% at only 5% of training indicates that even standard BF16 mixed-precision training may add non-trivial instability. If even mixed-precision BF16 can cause instabilities, the effects of using even fewer bits should be investigated further.
We make the following contributions in our work.
  • We analyze the effects of incremental bit reduction on training stability by simulating intermediate-bit representations of floating-point numbers. As observed in Kalamkar et al. [13], we confirm that reducing exponent bits has a greater impact on model performance than reducing mantissa bits, finding that removing even a single exponent bit on all layers precludes training altogether for Llama models.
@@
Figure 1: We show three cases of loss divergence on nanoGPT when using the same configurations as the default run except for the random seed. The blue lines indicate the average losses obtained for eight training runs that did not diverge. Of the 188 random seeds that were tested, 18 were found to diverge. As full pretraining requires over 4 days on a single node with 8 A100 GPUs, even for BF16, we perform early stopping at 30K steps, or 5% of the original training runs, requiring approximately 4 hours for a BF16 run and 8 hours for a TF32 run per A100 node with 8 GPUs. Because we only use 5% of the training data, we suspect that the measured divergence rate of approximately 10% underestimates the true rate of training loss divergence.
@@ 2.1 Training stability
Analyzing training instability in LLM pretraining directly is impractical due to the massive costs involved. Instead, smaller language models must be used as proxies. Wortsman et al. [18] explore the robustness of smaller language models as a proxy for LLM pretraining instability. They find that small models using a larger learning rate show similar instability patterns as larger models, showing unstable behavior, such as the growth of attention layer logits and the divergence of the output logits from the log probabilities, can be reproduced in smaller Transformers. They also explore both the causes of numerical instability in LLM training and mitigating strategies such as applying query/key (QK) normalization [19].
Keskar et al. [20] explores the sharpness of loss landscapes during large-batch neural network training, finding that larger batch sizes prevent the model from reaching flat regions of the loss landscape and causing performance degradation due to the inability to escape local minima. Of most significance to our work, they propose a metric for calculating loss landscape sharpness, which we adapt for LLMs.
@@ 2.2 Reduced-precision processors
To improve throughput on computationally intensive matrix multiplication tasks, recently developed processors have been equipped with specialized hardware units such as systolic arrays for TPUs [21] and tensor cores [22], which serve a similar purpose, for NVIDIA GPUs. These features can improve throughput by an order of magnitude. For example, for BF16, the maximum throughput on tensor cores is 312 TFLOPS compared to 39 TFLOPS using CUDA cores [22].
However, the number of multiplexer circuits required for the barrel shifter of an $n$-bit floating-point unit is $n\log_{2}n$ [23], which incentivizes the use of smaller floating-point representations. As a result, many mixed-precision techniques perform computationally intensive matrix multiplication tasks in BF16 while preserving sensitive portions of the model, such as the weights and residual path activations, in FP32. Alternatively, Henry et al. [24] developed a technique to approximate FP32 matrix multiplication using only BF16 by representing a single FP32 value as three BF16 values, which could accelerate FP32 matrix multiplication without requiring FP32 circuits.
@@ 2.3 Hybrid FP8
The adoption of the hybrid E5M2/E4M3 formats for neural networks [16] in recent generations of processors, such as the NVIDIA Hopper series of GPUs and the Intel Gaudi v2, has spurred interest in stable FP8 training. The hybrid FP8 format, where E4M3 is used for the forward pass for its greater resolution and E5M2 is used for the backward pass due to its greater expressive range, was first proposed by Wang et al. [8] as a means to accelerate both the training and inference stages of neural networks.
Sun et al. [15] built on this work to propose various techniques for stabilizing training, such as stochastic rounding, chunk-based accumulation, and Kahan summation during the optimizer update. However, the number of techniques that can be used in practice is limited by whether the technique in question can be applied without slowing the computation. As current NVIDIA GPUs do not support these techniques natively, the overhead caused by the software-based implementations cancels out any gains from the reduced precision.
@@ 2.4 MS-AMP
Introduced in Peng et al. [10], MS-AMP is an automatic mixed-precision package to utilize FP8 that offers multiple optimization levels to allow for differing model sensitivities when applying reduced precision to the computations and communications of neural network training. Our experiments use the O1 optimization level of MS-AMP, which performs GEMM computations in FP8 while reducing the weights to half-precision and uses FP8 states for the weight gradients and all-reduce communications. MS-AMP offers additional optimizations for the optimizer buffer in level O2 optimization and for distributed communication in level O3 optimization, but we choose to use only the most basic optimization scheme so as to verify the effects of the least invasive modifications.
@@ 3 Methods
We seek to answer whether sub-BF16 training is competitive with standard mixed-precision BF16 training from a cost-effectiveness point of view. In order to be cost-effective, reduced-precision training schemes must have minimal increases in training instability and changes to hyperparameters. To better analyze the effect of reduced precision on training stability, we aim to quantify the effects of gradually reducing floating-point representations bit by bit for both the exponent and the mantissa. Hopefully, analyzing the intermediate bit representations will better illuminate the interaction between bit width and training stability.
Our intermediate-bit floating-point experiments use the TinyLlama [25] repository, which implements the widely used Llama [5, 6] architecture. TinyLlama is an open-source implementation of Llama pretraining that uses a mix of the SlimPajama [26] and StarCoder [27] datasets. It also includes performance optimizations such as Flash Attention v2 [28]. We use the default learning rate of $lr=4e-4$, global batch size 512, and the same learning rate schedule as in the original code. The 120M models use a sequence length of 2048, while the 7B models use a sequence length of 4096.
@@ 3.1 Sharpness metric
Figure 2: Loss landscape diagrams for Llama 120M E8M3 at 5K steps (left) and 10K steps (right). Even during loss divergence, the loss landscape visualized using the method in [29] appears smooth, motivating our introduction of a new loss landscape sharpness metric. The value of the validation loss is included at each point of the loss landscape.
To better investigate the state of the model when loss divergence occurs, we attempted to visualize the loss landscape of the Llama models with the method proposed by Li et al. [29] in Figure 2. However, we found that even when the model is clearly in the process of loss divergence, the generated visualizations do not show signs of sharpness. Because of this, we propose a modified loss landscape sharpness metric that is more suitable for autoregressive models based on the one proposed in Keskar et al. [20]. The main difference with our version is that we use the logit of the last token instead of the model input for the calculation. This is because adding noise to the embeddings in a language model has different implications compared to adding noise to input images in a computer vision model. Instead of searching the input space as in Keskar et al. [20], we apply the search algorithm to the logit space of the last token. Searching the logit space has the additional advantage that the forward pass of the model need only be performed once for each measurement, greatly reducing the computational cost. The logit of the last token is chosen because it is not computationally feasible to perform an optimization on the entire output space. Also, due to the autoregressive character of decoder-only generative language models, the last token is the only one to receive inputs from all other tokens.
@@
The proposed metric can best be thought of as the relative magnitude of the largest loss spike on the logit within the provided bounds. The bounds are set to be the logit magnitudes plus one multiplied by $\epsilon$. The largest spike in the vicinity of the logits is found using the L-BFGS-B algorithm [30], using the SciPy [31] implementation with the output logit set as the starting point of the search. We use $\epsilon=5e\mathrm{-4}$ for all our experiments. Also, loss sharpness is measured with validation data that has not been used for training. In addition, we do not apply random projection as in Keskar et al. [20] to reduce the stochasticity of the measurement.

### v2

# To FP8 and Back Again: Quantifying Reduced Precision Effects on LLM Training Stability

@@ Abstract
The massive computational costs associated with large language model (LLM) pretraining have spurred great interest in reduced-precision floating-point representations to accelerate the process. As a result, the BrainFloat16 (BF16) precision has become the de facto standard for LLM training, with hardware support included in recent generations of accelerators. This trend has gone even further in the latest processors, where FP8 has recently been introduced. However, prior experience with FP16, which was found to be less stable than BF16, raises concerns as to whether FP8, with even fewer bits than FP16, can be a cost-effective option for LLM training. We argue that reduced-precision training schemes must have similar training stability and hyperparameter sensitivities to their higher-precision counterparts in order to be cost-effective. However, we find that currently available methods for FP8 training are not robust enough to allow their use as economical replacements. This prompts us to investigate the stability of reduced-precision LLM training in terms of robustness across random seeds, learning rates, and datasets. To this end, we propose new evaluation techniques and a new metric for quantifying loss landscape sharpness in autoregressive language models. By simulating incremental bit reductions in floating-point representations, we analyze the relationship between representational power and training stability with the intent of aiding future research into the field.
LLM, FP8, Training, Stability, Cost
@@ 1 Introduction
Conversational large language models (LLMs), such as ChatGPT (OpenAI, 2024), Gemini (Team, 2024a, b), Claude (Anthropic, 2024), DeepSeek (DeepSeek-AI, 2024) and HyperCLOVA (Yoo et al., 2024), have captured the imagination of both academics and the public at large with their ability to communicate fluently with humans in natural language. However, these models require unprecedented amounts of computation to train, which has engendered interest in methods to improve their training efficiency.
A popular method of improving computational performance is to reduce the bit count of the floating-point representations used for training (Wang et al., 2018; Sun et al., 2020; Peng et al., 2023). Because reading memory is the main bottleneck in modern processors, a problem known as the “memory wall” (Wulf & McKee, 1995; Kim et al., 2023), reducing the number of bits that each floating-point number uses can accelerate the computation in proportion to the amount of memory reduced. For example, in processors that support it, computations in BrainFloat16 (BF16) (Kalamkar et al., 2019) can have double the maximum throughput of single precision FP32. Furthermore, the FP32 data type, the highest precision data type used in deep learning, has only half the bits of FP64, the most widely used floating-point data type in scientific computing. The current best practice for LLM training is to use BF16 for most of the LLM training computation, with some sensitive portions, such as layer normalization (Ba et al., 2016), carried out in FP32.
As a natural extension of this development, 8-bit floating-point (FP8) (Wang et al., 2018; Sun et al., 2019; Micikevicius et al., 2022; Peng et al., 2023) and even 4-bit floating-point (Sun et al., 2020) data formats have been proposed to accelerate training even further. However, the naïve application of FP8 to LLM training is unstable and requires additional techniques to become viable (Shen et al., 2024). While several methods have been proposed to stabilize training LLMs with FP8, relatively little attention has been paid to quantifying the decrease in stability compared to mixed-precision BF16 training.
Cost reduction is the motivation behind the use of FP8 and other reduced-precision training schemes as there is no intrinsic benefit in training models using smaller data types. Therefore, our concern is not whether LLM pretraining with FP8 is possible but whether it is profitable. For cost savings to be realized, the time per training step must be reduced while the number of training steps is kept to a similar number. Training stability is thus a crucial factor for cost-effective LLM training, considering that additional hyperparameter searches and restarts from training failures can outweigh any gains from raw compute acceleration.
@@
We motivate our line of inquiry with some surprising findings from experiments on the nanoGPT (Karpathy, 2022) codebase, an open-source implementation of GPT-2 pretraining, where we found that even the current best practice of mixed-precision BF16 can introduce training instabilities. When we compared BF16 and TensorFloat32 (TF32) runs, where we ran training for 5% of the original configuration, we found that the BF16 models diverged for 18 of 188 runs, or approximately 10% of all cases, despite using the same configurations as the default run. In contrast, no cases of loss divergence were found for the 70 TF32 models trained using different random seeds. We use the TF32 data type because NVIDIA GPUs do not offer FP32 tensor cores.
This is surprising in light of the fact that most recent LLMs are trained with mixed precision BF16 without a comparison with training on TF32, which has three additional mantissa bits. However, a loss divergence rate of approximately 10% at only 5% of training indicates that even standard BF16 mixed-precision training may add non-trivial instability. If even mixed-precision BF16 can cause instabilities, the effects of using even fewer bits should be investigated further.
We make the following contributions in our work:
  • We analyze the narrower hyperparameter space that emerges from reducing precision by clipping mantissa bits to simulate intermediate-bit floating-point representations and find greater instability in the model when exposed to higher learning rates or “dirtier” data.
@@
To prevent misunderstanding, we also clarify on points we do not make. We do not argue that LLM pretraining with FP8 is impractical, only that there must be a deeper investigation into whether proposed methods are economical. While various techniques for FP8 training (Fishman et al., 2024; Peng et al., 2023; Xi et al., 2025) and even FP4 training (Wang et al., 2025) have been proposed, what remains unclear is whether these methods results in a cost reduction when considering factors other than raw compute speed.
During training, a single loss divergence makes redundant the GPU time from the last checkpoint up to the idle period during the restart, possibly canceling out any gains from the accelerated computation. Also, during actual FP8 training, we cannot know what the training curve would be had BF16 been used instead. If the model requires additional steps to converge to the same point, this would further increase the cost. Even if training works reasonably well on standardized datasets and configurations, practitioners wish to know whether these findings are generalizable. We seek not to discredit any particular method but to show how a new training method can prove itself robust and trustworthy.
Figure 1: We show an example case of loss divergence on nanoGPT when using the same configurations as the default run except for the random seed. The blue lines indicate the average losses obtained for eight training runs that did not diverge. Of the 188 random seeds that were tested, 18 were found to diverge. As full pretraining requires over 4 days on a single node with 8 A100 GPUs, even for BF16, we perform early stopping at 30K steps, or 5% of the original training steps, requiring approximately 4 hours for a BF16 run and 8 hours for a TF32 run per A100 node with 8 GPUs. Because we only run 5% of the original training, we suspect that the measured divergence rate of approximately 10% underestimates the true rate of training loss divergence.
@@ 2 Related work
2.1 DeepSeek V3
As we were preparing this work, DeepSeek V3 (DeepSeek-AI, 2024) has gained immense popularity as a open-source SOTA model that is capable of rivaling closed-source competitors such as ChatGPT o1. Notably, it proposes a sophisticated FP8 quantization scheme with fine-grained FP8 dynamic scaling with a group size of 128 and accumulation to FP32 for every 4 WGMMA instructions to compensate for the 14-bit accumulation of NVIDIA FP8 tensor cores. This raises a question. Had previously proposed FP8 training schemes been sufficiently stable to train LLMs cost effectively, it would not be necessary to write customized kernels. For example, in none of the previous works (Xi et al., 2025; Fishman et al., 2024; Peng et al., 2023; NVIDIA, 2023) on FP8 training was a scaling factor granularity of 128 proposed, with per-row scaling being the finest granularity used. Furthermore, prior to DeepSeek V3, the backward pass was implemented using E5M2, with it being the first work that we are aware of to apply E4M3 to the backward pass as well as the forward pass.
We take the unconventional FP8 implementation in the work as confirmation of our claim that previous attempts at FP8 training had not gained economical viability. Furthermore, we argue that new implementations of FP8 should undergo rigorous analysis on whether they truly are equivalent to BF16 training, and not overfit to the architectures and hyperparameters used for well-known tasks.
@@ 2.2 Training stability
Analyzing training instability in LLM pretraining directly is impractical due to the massive costs involved. Instead, smaller language models must be used as proxies. Wortsman et al. (2024) explore the robustness of smaller language models as a proxy for LLM pretraining instability. They find that small models using a larger learning rate show similar instability patterns as larger models, such as the growth of attention layer logits and the divergence of the output logits from the log probabilities. They also explore both the causes of numerical instability in LLM training and mitigating strategies such as applying query/key (QK) normalization (Dehghani et al., 2023).
Keskar et al. (2017) explore the sharpness of loss landscapes during large-batch neural network training, finding that larger batch sizes prevent the model from reaching flat regions of the loss landscape and causing performance degradation due to the inability to escape local minima. Of most significance to our work, they propose a metric for calculating loss landscape sharpness, which we adapt for LLMs as a proxy for training instability.
Fishman et al. (2024) go further, discovering that Llama 7B models trained in FP8 begin to diverge after 200B tokens of training, supporting our claim that reduced-precision methods induce hidden instabilities that may emerge only later in training. They identify SwiGLU (Shazeer, 2020) as a source of massive activations (Sun et al., 2024) and apply additional scaling factors to reduce the instability. However, it remains unclear whether the proposed method is equivalent to BF16 training or are simply stable enough for the experiments involved.
@@ 2.3 Reduced-precision processors
To improve throughput on computationally intensive matrix multiplication tasks, recently developed processors have been equipped with specialized hardware units such as systolic arrays for TPUs (Jouppi et al., 2017) and tensor cores (NVIDIA, 2020) for NVIDIA GPUs. These processors can improve throughput by an order of magnitude. For example, on the H100, the peak BF16 GEMM throughput on tensor cores is 989.4 TFLOPS, compared to 133.8 TFLOPS on CUDA cores (NVIDIA, 2022).
However, the number of multiplexer circuits required for the barrel shifter of an $n$-bit floating-point unit is $n\log_{2}n$ (Kroening & Strichman, 2008), which incentivizes using smaller floating-point representations. As a result, many mixed-precision techniques perform computationally intensive matrix multiplication in BF16 while preserving sensitive portions of the model, such as the master weights and residual path activations, in FP32.
@@ 2.4 Hybrid FP8
The adoption of the hybrid E4M3/E5M2 formats for neural networks (Micikevicius et al., 2022) in recent generations of processors, such as the NVIDIA H100 and the Intel Gaudi v2, has spurred interest in stable FP8 training. The hybrid FP8 format, where E4M3 is used for the forward pass for its greater resolution, and E5M2 is used for the backward pass for its greater range, was first proposed by Wang et al. (2018) as a means to accelerate neural networks.
Sun et al. (2019) built on this work to propose various techniques for stabilizing training, such as stochastic rounding, chunk-based accumulation, and Kahan summation during the optimizer update. However, the number of techniques that can be used in practice is limited by whether the technique in question can be applied without slowing the computation. As currently available NVIDIA GPUs, by far the accelerators with the greatest adoption, do not support these techniques natively, the overhead caused by the software-based implementations cancels out any gains from the reduced precision computations.
@@ 2.5 MS-AMP
Introduced in Peng et al. (2023), MS-AMP is an automatic mixed-precision package to utilize FP8 that offers multiple optimization levels to allow for differing model sensitivities when applying reduced precision to the computations and communications of neural network training. Our experiments use the O1 optimization level of MS-AMP, which performs GEMM computations in FP8 while reducing the weights to half-precision and uses FP8 states for the weight gradients and all-reduce communications. MS-AMP offers additional optimizations for the optimizer buffer in level O2 optimization and for distributed communication in level O3 optimization, but we use only the most basic scheme so as to verify the effects of the least invasive modifications.
@@ 3 Methods
We seek to answer whether sub-BF16 training is competitive with standard mixed-precision BF16 training from a cost-effectiveness point of view. To be cost-effective, reduced-precision training schemes must have minimal increases in training instability and changes to hyperparameters. To better analyze the effect of reduced precision on training stability, we aim to quantify the effects of gradually reducing floating-point representations bit by bit. Hopefully, analyzing the intermediate bit representations will better illuminate the interaction between bit width and training stability.
Our intermediate-bit floating-point experiments use the TinyLlama (Zhang et al., 2024) repository, which implements the widely used Llama (Touvron et al., 2023a, b) architecture. TinyLlama is an open-source implementation of Llama pretraining that uses a mix of the SlimPajama (Soboleva et al., 2023) and StarCoder (Li et al., 2023) datasets for training. It also includes performance optimizations such as Flash Attention v2 (Dao, 2023). We use the default learning rate of $lr=4e-4$, global batch size 512, and the same learning rate schedule as in the original code. The 120M models use a sequence length of 2048, while the 7B models use a sequence length of 4096.
@@ 3.1 Sharpness metric
Figure 2: Loss landscape diagrams for Llama 120M E8M3 at 5K steps (left) and 10K steps (right). Even during loss divergence, the loss landscape visualized using the method in Li et al. (2018) appears smooth, motivating our introduction of a new loss landscape sharpness metric. The validation loss is shown for each point of the loss landscape.
{minted}
python def forward(x, w): # x: input, w: weight, out: output save_for_backward(x, w) masked_x = reduce_precision(x) masked_w = reduce_precision(w) out = F.linear(masked_x, masked_w) masked_out = reduce_precision(out) return masked_out
Figure 3: PyTorch-like pseudocode for the forward pass.
To better investigate the model state when loss divergence occurs, we attempted to visualize the loss landscape of the Llama models with the method proposed by Li et al. (2018). However, as shown in Figure 2, we found that even when the model is clearly in the process of loss divergence, the generated visualizations remain smooth. We emphasize that we do not seek the loss landscape sharpness per se, but a proxy to provide a quantitative measurement of the underlying training instability.
Because of this, we propose an alternative loss landscape sharpness metric that is more suitable for autoregressive models, based on the one proposed in Keskar et al. (2017). We empirically confirm that it is a useful indicator of training instability in the following sections. The main difference between the original metric and our version is that we use the logit of the last token instead of the model input for the calculation. This is because adding noise to the embeddings in a language model has different implications compared to adding noise to input images in a vision model. We also do not apply random projection as in Keskar et al. (2017) to reduce the stochasticity of the measurement.
Instead of searching the input space as in Keskar et al. (2017), we apply the search algorithm to the logit space of the last token. Searching the logit space has the additional advantage that the forward pass of the model need only be performed once for each measurement, significantly reducing the computational cost. The logit of the last token was chosen because it is not computationally feasible to optimize for the entire output space. Also, due to the autoregressive character of decoder-only Transformer models (Vaswani et al., 2017; Brown et al., 2020), the last token is the only one to receive inputs from all other tokens.
@@
The proposed metric can best be thought of as the relative magnitude of the largest loss spike on the logit within the provided bounds. The bounds are set to be the logit magnitudes plus one multiplied by $\epsilon$. The largest spike in the vicinity of the logits is found using the L-BFGS-B algorithm (Liu & Nocedal, 1989), using the SciPy (Virtanen et al., 2020) implementation with the output logit set as the starting point of the search. We set $\epsilon=5e\mathrm{-4}$ for all our experiments following Keskar et al. (2017). However, a hyperparameter sweep on $\epsilon$ shows that the general trend is unaffected by the value of $\epsilon$, only the sharpness value magnitudes. The results are shown in Table To FP8 and Back Again: Quantifying Reduced Precision Effects on LLM Training Stability of the Appendix.